OpenStudy (anonymous):

Solve |2 - 3x| = 7. {-5/3, 3} {-15, 3} {-3, 5/3} {-3, 15}

OpenStudy (anonymous):

|2-3x| = 7 2-3x= \[\pm\]7 2-3x=+7 -3x=5 x=-5/3 2-3x =-7 -3x= -9 x=3

OpenStudy (anonymous):

3, 5/3

OpenStudy (anonymous):

* -

hero (hero):

|2 - 3x| = 7 2 - 3x = -7 2 - 3x = 7 2 +7 = 3x 2 - 7 = 3x 9 = 3x -5 = 3x 3 = x -5/3 = x Solution: -5/3, 3